1.)What is the payment on a 60-month, $10,000 car loan with APR of 6.66%? Round to the penny.

2.) In the first payment on a 60-month, $10,000 car loan with APR of 3.65%, how much pays off the principal? Round to the penny.

3.)After making 28 payments on a 60-month, $10,000 car loan with APR of 6%, what is the outstanding balance on the loan? Round to the penny.
Answer #1

1.

Monthly payment is:

Cm Loan amount = Monthly payment 0.0666 11+- 12 0.0666 10,000 = Monthly payment 12 10,000 = Monthly payment x 50.9134402 196.

2.

12 0.0365 1-11+- 10,000 = Monthly payment | 0.0365 12 10,000 = Monthly payment * 54.76751995 182.59 = Monthly payment

The interest part in first payment would be $10,000*(0.0365/12) = $30.4167

Thus, principal amount must be ($182.59-$30.4167) = $152.17.

3.

10,000 = Monthly paymentx 0.06 10,000 = Monthly payment x 51.72556075 193.33= Monthly payment

Outstanding balance is:

11+ Present value = Monthly payments 1-(1+0:06) = 193.33% 193.3280153 12 0.06 12 = 193.33x 29.50328355 = 5,703.81